A Package to Set Margins to Full Page

Patrick W. Daly

This paper describes package fullpage version 1.1 from 1999/02/23

It is part of the preprint collection of packages

Summary

The stripped version of this file contains the following brief description:

 % Sets all 4 margins to be either 1 inch or 1.5 cm, and specifies
 %  the page style
 % Call as    \usepackage[options]{fullpage}
 % Options  =   in   (default) for 1 inch
 %              cm             for 1.5 cm
 %              plain (default) sets \pagestyle{plain}
 %              empty           sets \pagestyle{empty}
 %              headings        sets \pagestyle{headings}
 %              myheadings      sets \pagestyle{myheadings}

1 Purpose

This package is called fullpage, and it sets the text height and width and the margins such that the page is almost full. It only leaves a uniform margin on all four sides.

To do this, it needs to know how big the full page is, whether US letter size (8\frac{1}{2} × 11′′) or A4 (21 × 29.7 cm), or some other size. This is given in the parameters \paperwidth and \papersize that are set by the appropriate option in the

\documentclass command.

The package must also know which page style is being employed. This is because it must leave space for possible headers and footers, if they are present. If not, it may use that space for text.

2 Invoking the Package

The package is called after the \documentclass command as e.g.,

     \documentclass[a4paper,landscape]{article}
     \usepackage[options]{fullpage}

where possible options for fullpage are in (default) sets the margins to one inch; cm sets the margins to 1.5 cm (one centimeter is really too little); plain (default) selects the plain page style, i.e., with no headers but

only a footer; empty for neither headers nor footers; headings for both header and footers; myheadings also for both headers and footers.

For the last 4 options, the corresponding \pagestyle declaration is executed, so that it is not necessary to give it again.